### v2

@@ 3.3 Training and Inference
Based on Eq. 2, the image classification loss is formulated as:
@@
In addition to $\mathcal{L}_{ce}$, a matching loss is needed to pull the matched top-$K_{3}$ keys $\mathcal{K}_{j}$ closer to the image embedding $\mathbf{z}_{j}$, so that the keys learn rich concepts from the samples. We use cosine distance as the matching loss. However, other similar criteria, like Euclidean distance, can also be used as a matching loss. We empiricaly observe that cosine distance as the matching loss yields the best performance. Therefore The matching loss adopted to optimize the keys is defined as:
@@
Finally, in order to make the learned prompts more semantically diverse, we adopt a third loss to orthogonalize the embeddings of different prompts to increase the diversity of the prompts:
@@
The keys are optimized by $\mathcal{L}_{ma}$, and the prompts by $\mathcal{L}_{ce}$, $\mathcal{L}_{or}$ and $\mathcal{L}_{cc}$.
@@
Once visual and textual concepts (keys and values) are learned through training, they can be shipped with CLIP for downstream tasks with a standard zero-shot CLIP inference setup. As shown in Figure 3, we first generate the visual features $f_{vt}$ of the test image using visual Encoder $E_{v}$, then we use $f_{vt}$ to choose top-$K_{3}$ similar visual concepts(keys) by cosine similarity, next, we concatenate the corresponding prompts(values) of the keys to obtain the conceptual prompt, which is fused with each given name to produce conceptual prompt text features $\{f_{tt}^{i}\}_{i=1}^{C}$. Finally, zero-shot inference is performed with the conceptual prompted text features and the input image feature $f_{vt}$ to produce classification scores on the test images.
@@ 4 Experiments
In this section, We present a thorough evaluation of our method, including experimental results, performance comparisons with state-of-the-art methods, and ablation studies.
@@ 4.1 Experimental Setup
  • Base to Novel Class generalization. We evaluate the generalization capability of our method in zero-shot scenarios within a dataset. The dataset is evenly divided into base and novel classes. We train our model with few-shot images on the base classes and then evaluate on both base classes and unseen novel classes.
@@
  • Domain Generalization. We evaluate our model’s performance on out-of-distribution generalization. Similarly, we assess our model trained on ImageNet directly on four ImageNet variants that contain the same classes but from different distributions
@@
Table 1: Comparison with state-of-the-art methods on base-to-novel generalization. The best accuracies are bolded and the second best results are underlined. HM indicates the harmonic mean.
@@
To ensure a fair comparison, we employ the ViT-B/16 CLIP model across all three benchmark tasks. For base-to-novel generalization, we train our model with 16-shot images on base classes and then evaluate both base classes and novel classes. For cross-dataset evaluation and domain generalization, both use the model trained with 16-shot ImageNet and test on each target dataset. Throughout the training, we keep both the visual and textual encoders fixed. Our data preprocessing follows CLIP’s protocol, including resizing and random cropping operations, among others. For the base-to-novel generation task, we train for 30 epochs on ImageNet and 20 epochs on other datasets. $K_{1}$ and $K_{2}$ are set to 3 and 10 respectively. We set prompt length $M$ to 8, the number of concepts in the bank $N$ to 100, and the number of selected concepts $K_{3}$ to 4. Training utilizes a batch size of 8 with an initial learning rate of $10^{-3}$. We utilize the AdamW optimizer alongside a cosine annealing scheduler. Results are averaged over 3 runs for all methods.
@@ 4.2 Base-to-Novel Generalization
In this section, We compare our proposed method with seven baselines: zero-shot CLIP [26], CoOp [42], CoCoOp [41], MaPLe [17], KgCoOp [36], CoPrompt [28] and CPL [40]. The experimental results regarding base-to-novel generalization across 11 datasets with 16-shot samples are presented in Table 1. We have highlighted the best results in bold and marked the improvement over the second-best performance in blue. As we can see from Table 1(a), the average of all 11 datasets shows that our method outperforms all the baselines by a large margin for both base and novel classes. In comparison to CoOp and CoCoOp, which are the pioneering prompt learning methods, the performance gain of our method even reached +11% and +6.9% respectively. Our method outperforms the previous state-of-the-art (CPL) by +2.28% on novel classes and +1.62% on the harmonic mean (HM). These results demonstrate the strong zero-shot generalization capability of our proposed method. Also, our method outperforms CPL on base classes by +0.84%, which shows a strong few-shot learning capability.
@@ 4.3 Cross-Dataset Evaluation
The comparison results with CoOp, CocoOp, MaPLe, CoPrompt, and CPL are presented in Table 2. Overall, our CoCoLe method marks the best performance on both source and target datasets with a target average of 68.91%, and outperforms CPL by 0.84%. Notably, we obtain the highest improvement of 2.26% over CPL on UCF101, an action image dataset whose fundamentals are distinctive from ImageNet. This suggests that conceptual codebook learning in our method facilitates better generalization.
Table 2: Comparison of our method with existing approaches on cross-dataset evaluation. Overall, our method demonstrates superior generalization capabilities with the highest average accuracy on 10 datasets.
@@ 4.4 Domain Generalization
In Table 3, we provide the classification accuracy across the source domain and target domains, as well as the average accuracy within target domains. Our approach consistently outperforms all baselines on both source and target datasets, achieving a new state-of-the-art average accuracy of 61.85% on this task. Notably, our method beats CPL [40] by +1.51% on ImageNet-R. This indicates the remarkable robustness of our model against distribution shifts.
[…]
@@ 4.5 Ablation Studies
To systematically evaluate our proposed method, we provide an empirical analysis of our design choices and illustrate the effects of different components of our method in this section. If not otherwise specified, our experiments are conducted on the ImageNet for the Base-to-Novel Task, and we report the harmonic mean.
@@
In Tab. 4, $\mathcal{L}_{ma}$ is the matching loss adopted to optimize the keys, $\mathcal{L}_{cc}$ is for consistency constraint, $\mathcal{L}_{or}$ is used to orthogonalize the embeddings of different prompts. We present an ablation study by removing different components of the proposed method to understand the importance of each of them. For inference, in the first row of the table, we present the final performance of CoCoLe, which has a harmonic mean of 76.84%. In the first ablation experiment, we eliminate the $\mathcal{L}_{or}$ from CoCoLe, leading to a performance drop of 0.46%. This highlights the importance of $\mathcal{L}_{or}$ for CoCoLe. Next, we remove $\mathcal{L}_{cc}$, effectively enforcing consistency between the learnable conceptual prompts and handcrafted concept-based prompts. This results in a 1.98% drop in performance, suggesting the importance of consistency regularization. Finally, we remove $\mathcal{L}_{ma}$, which leads to a performance drop of 2.39%. This shows that the learnable conceptual codebook plays a crucial role in CoCoLe. In general, the overall results highlight the significant contribution of all components to enhanced performance.
Table 4: The ablation study on each component of CoCoLe. The experiments are conducted on ImageNet. We report the harmonic mean accuracy of the base-to-novel task.
@@
Table 5: Ablation study of the length of prompts $M$, the size of conceptual codebook $N$, the number $K_{3}$ of visual concepts for training at a time.
@@
As shown in Tab. 6, our proposed CoCoLe exhibits a significant performance advantage over other methods. Though our CoCoLe requires more training time compared to CoOp, it still outperforms CoOp by 11%. Furthermore, when compared to CoCoOp and CPL, our method achieves substantial performance gains with less time. This demonstrates the efficiency and effectiveness of our method.
@@
As shown in Fig. 1, in order to confirm that various prompts indeed capture distinct image concepts, we employ Grad-CAM [30] to visually represent the image contents associated with different prompts. Observing Fig. 1(a), it’s apparent that various prompts highlight distinct regions within the same image, showcasing the diversity of the acquired prompts. For instance, different prompts applied to the Koala image emphasize different areas such as the head, claws, and tree, illustrating the versatility of the learned prompts. To assess whether the learned prompts indeed encapsulate image concepts with higher-level semantics, we visualize the content of specific prompts ($\textbf{P}_{3}$, $\textbf{P}_{23}$, $\textbf{P}_{48}$) across different images in Fig. 1(b). It’s evident that $\textbf{P}_{3}$ mainly captures the concept "wheels," $\textbf{P}_{23}$ encapsulates the concept "grass," while $\textbf{P}_{48}$ focuses on the "screen" of the devices. This highlights the effectiveness of prompts in learning key concepts that can be generalized across images, thereby enhancing performance in generalization tasks. In the Supplemental Materials, we provide additional details of the proposed method and experimental results.
@@ 5 Conclusion
To address the challenge of improving the generalization capability of VLMs while fine-tuning them on downstream tasks, we propose Conceptual Codebook Learning (CoCoLe). The learned conceptual codebook consists of visual concepts as keys and conceptual prompts as values, which serves as a link between the image encoder’s outputs and the text encoder’s inputs. Additionally, we incorporate a handcrafted concept cache as a regularization to alleviate the overfitting issues in low-shot scenarios. Extensive experimental results demonstrate that our CoCoLe method remarkably outperforms the existing state-of-the-art methods.

### v3

@@ 3.3 Training and Inference
According to Equation 1, the loss for image classification is expressed as:
@@
Besides $\mathcal{L}_{ce}$, we require a matching loss that brings the matched top-$K_{3}$ keys $\mathcal{K}_{j}$ closer to the image embedding $\mathbf{z}_{j}$, facilitating the keys to learn diverse concepts from the samples. We employ cosine distance as our matching loss. Nevertheless, alternative metrics such as Euclidean distance can also serve as a viable matching loss. Through empirical observation, we find that cosine distance typically delivers optimal performance as a matching loss. Therefore, The matching loss adopted to optimize the keys is defined as:
@@
Lastly, to enhance the semantic diversity of the learned prompts, we introduce an additional loss that orthogonalizes the embeddings of different prompts, thereby boosting prompt diversity.
@@
The keys in the conceptual codebook are optimized by $\mathcal{L}_{ma}$, while the prompts are optimized through $\mathcal{L}_{ce}$, $\mathcal{L}_{or}$ and $\mathcal{L}_{cc}$.
@@
Once visual and textual concepts (keys and values) are learned through training, they can be shipped with CLIP for downstream tasks with a standard zero-shot CLIP inference setup. As shown in Figure 3, we first generate the visual features $f_{vt}$ of the test image using visual Encoder $E_{v}$, then we use $f_{vt}$ to choose top-$K_{3}$ similar visual concepts(keys) by cosine similarity, next, we concatenate the corresponding prompts(values) of the keys to obtain the conceptual prompt, which is fused with each given name to produce conceptual prompt text features $\{f_{tt}^{i}\}_{i=1}^{C}$. Finally, the zero-shot inference is performed with the conceptual prompted text features and the input image feature $f_{vt}$ to produce classification scores on the test images.
@@ 4.1 Experimental Setup
  • Base to Novel Class generalization. We evaluate the generalization capability of our method in zero-shot scenarios within a dataset. The dataset is evenly divided into base and novel classes. We train our model with few-shot images on the base classes and then do the evaluation on both base classes and unseen novel classes.
@@
  • Domain Generalization. We assess the performance of our model on out-of-distribution generalization. Likewise, we evaluate our model trained on ImageNet directly on four variants of ImageNet, each containing the same classes but from different distributions.
@@
Table 1: Comparison with existing methods on base-to-novel generalization. The top accuracies are highlighted in bold, with the second-best results underlined. The harmonic mean is denoted by HM.
@@
To ensure a fair comparison, we employ the ViT-B/16 CLIP model across all three benchmark tasks. For base-to-novel generalization, we train our proposed CoCoLe with 16-shot images on base classes and subsequently evaluate it in both base classes and novel classes. For domain generalization and cross-dataset evaluation, we utilize the model trained with 16-shot ImageNet and test it on each target dataset. Throughout the training, we keep both the visual and textual encoders fixed. Our data preprocessing follows CLIP’s protocol, including resizing and random cropping operations, among others. For the base-to-novel generation task, we conduct training for 30 epochs on ImageNet and 20 epochs on the other datasets. $K_{1}$ and $K_{2}$ are set to 3 and 10 respectively. We set prompt length $M$ to 8, the size $N$ of the conceptual codebook to 100, and the number of selected concepts $K_{3}$ to 4. The training is performed with a batch size of 8 with an initial learning rate of $10^{-3}$. We utilize the AdamW optimizer alongside a cosine annealing scheduler.
@@ 4.2 Base-to-Novel Generalization
In this section, We compare our CoCoLe method with seven baselines: zero-shot CLIP [26], CoOp [42], CoCoOp [41], MaPLe [17], KgCoOp [36], CoPrompt [28] and CPL [40]. Table 1 shows the experimental results for base-to-novel generalization across 11 datasets using 16-shot samples. We have bolded the top results and indicated improvements over the second-best performance in blue. As we can see from Table 1(a), the average of all 11 datasets shows that our method outperforms all the baselines by a large margin for both base and novel classes. In comparison to CoOp and CoCoOp, which are the pioneering prompt learning methods, the performance gain of our method even reached +11% and +6.9% respectively. Our method outperforms the previous state-of-the-art (CPL) by +2.28% on novel classes and +1.62% on the harmonic mean (HM). These results demonstrate the strong zero-shot generalization capability of our proposed method. Also, our method outperforms CPL on base classes by +0.84%, which shows a strong few-shot learning capability.
@@ 4.3 Cross-Dataset Evaluation
Table 2 displays the comparison results with CoOp, CocoOp, MaPLe, CoPrompt, and CPL. Our CoCoLe approach achieves the top performance on both source and target datasets, averaging 68.91% on the target set, surpassing CPL by 0.84%. Notably, we observe the largest improvement of 2.26% over CPL on UCF101, an action image dataset with distinct characteristics from ImageNet. This underscores that our method’s conceptual codebook learning enhances generalization significantly.
Table 2: Comparison with state-of-the-art methods on cross-dataset evaluation.
@@ 4.4 Domain Generalization
We present the classification accuracy for both the source domain and target domains in Table 3, along with the average accuracy across the target domains. Our approach consistently outperforms all baselines on both source and target datasets, setting a new state-of-the-art average accuracy of 61.85% for this domain generalization task. Notably, our method beats CPL [40] by +1.51% on ImageNet-R. This highlights the exceptional robustness of our model against distribution shifts.
Table 3: Comparison with existing methods on domain generalization task. The top results are highlighted in bold with the second-best results underlined.
@@ 4.5 Ablation Studies
To rigorously assess our proposed approach, we conduct an empirical analysis of our design decisions and demonstrate the impact of various components in this section. Unless otherwise stated, our experiments are conducted on ImageNet for the Base-to-Novel Task, and we report the harmonic mean.
@@
In Tab. 4, $\mathcal{L}_{ma}$ is used as a matching loss to optimize the visual concepts in the conceptual codebook. $\mathcal{L}_{cc}$ is for consistency constraint, while $\mathcal{L}_{or}$ is employed to orthogonalize the textual features of different prompts. We conduct an ablation study by systematically removing various components of our proposed CoCoLe to assess their individual importance. For evaluation, the first row of the table showcases the overall performance of CoCoLe, achieving a harmonic mean of 76.84%. In the initial ablation experiment, we exclude $\mathcal{L}{or}$ from CoCoLe, resulting in a performance decrease of 0.46%. This underscores the significance of $\mathcal{L}{or}$ in CoCoLe. Next, we remove $\mathcal{L}_{cc}$, effectively enforcing consistency between the learnable conceptual prompts and handcrafted concept-based prompts. As a result, there is a decrease in performance by 1.98%, indicating the significance of the regularization strategy. Finally, we remove $\mathcal{L}_{ma}$, which leads to a performance drop of 2.39%. This shows that the learnable conceptual codebook plays a crucial role in CoCoLe. In general, the overall results highlight the significant contribution of all components to enhanced performance.
Table 4: The ablation study on each component of CoCoLe.
@@
Table 5: Ablation study of $M$, $N$, and $K_{3}$.
@@
As shown in Tab. 6, our proposed CoCoLe exhibits a significant performance advantage over other methods. Though our CoCoLe requires more training time compared to CoOp, it still outperforms CoOp by 11%. Furthermore, when compared to CoCoOp and CPL, our method achieves substantial performance gains with less time. This showcases the efficiency and effectiveness of our proposed CoCoLe.
@@
As shown in Fig. 1, in order to confirm that various prompts indeed capture distinct image concepts, we employ Grad-CAM [30] to visually represent the image contents associated with different prompts. Observing Fig. 1(a), it’s apparent that various prompts highlight distinct regions within the same image, showcasing the diversity of the acquired prompts. For instance, different prompts applied to the Koala image emphasize different areas such as the head, claws, and tree, illustrating the versatility of the learned prompts. To determine if the learned prompts indeed encapsulate higher-level semantics image concepts, we visualize the content of specific prompts ($\textbf{P}_{3}$, $\textbf{P}_{23}$, $\textbf{P}_{48}$) across different images in Fig. 1(b). It’s evident that $\textbf{P}_{3}$ mainly captures the concept "wheels," $\textbf{P}_{23}$ encapsulates the concept "grass," while $\textbf{P}_{48}$ focuses on the "screen" of the devices. This highlights the effectiveness of prompts in learning key concepts that can be generalized across images, thereby enhancing performance in generalization tasks. In the Supplemental Materials, we provide additional details of our proposed CoCoLe and experimental results.
@@ 5 Conclusion
To address the problem of enhancing the generalization capability of VLMs while adapting them to downstream tasks, we propose Conceptual Codebook Learning (CoCoLe). The learned conceptual codebook consists of visual concepts as keys and conceptual prompts as values, which serve as a link between the image encoder’s outputs and the text encoder’s inputs. Additionally, we incorporate a handcrafted concept cache as a regularization to alleviate the overfitting issues in low-shot scenarios. Extensive experimental results demonstrate that our CoCoLe method remarkably outperforms the existing state-of-the-art methods.
Acknowledgements
This research was supported by the National Natural Science Foundation of China (No. 62331014) and Project 2021JC02X103.